A 15-g sample of the candy is placed in a small aluminum container of mass 0.325 kg filled with oxygen. This container is placed in 1.75 kg of water in an aluminum calorimeter cup of mass 0.624 kg at an initial temperature of 15.0 ∘C. The oxygen-candy mixture in the small container (a "bomb calorimeter") is ignited, and the final temperature of the whole system is 53.5 ∘C.The value of specific heat for aluminum is 0.22 Cal/kg⋅C∘, and for water is 1.00 Cal/kg⋅C∘ Determine the Calorie content of 77 g of candy.
Answer #1

Step 1: Find calorie content released by burning of 15 gm candy:

Q = Q1 + Q2

Q1 = Heat released by aluminum = m1*C1*dT

m1 = mass of aluminum container + calorimeter = 0.325 + 0.624 = 0.949 kg

C1 = Specific heat capacity of Aluminum = 0.22 Cal/lg.C

dT = 53.5 - 15 = 38.5 C

Q2 = Heat released by water = m2*C2*dT

m2 = mass of water = 1.75 kg

C2 = specific heat capacity of water = 1.00 Cal/kg-C

So,

Q = 0.949*0.22*38.5 + 1.75*1.00*38.5 = 75.413 cal

Now Since above calorie content is released by 15 gm candy, So

Calorie content in 77 gm candy = 75.413*77/15

Calorie content in 77 gm candy = 387.12 cal

Let me know if you've any query.